How many minutes does it take to travel 10 miles at 80 miles per hour?
Answer by stanbon(75887)

How many minutes does it take to travel 10 miles at 80 miles per hour?
-----
Equation:
d/r = time
----
10 miles / 80 miles/hr = (1/8)hr = (1/8)*(60 min) = 7.5 minutes
----------------------
Cheers,
Stan H.
--------------------